OpenStudy (anonymous):

A=1/2h(a+b) solve for a

OpenStudy (anonymous):

So, we have: \[ \begin{align} A&=\frac{1}{2}h(a+b)\implies\\ 2A&=h(a+b)\implies\\ \frac{2A}{h}&=a+b\implies\\ \frac{2A}{h}-b&=a \end{align} \]